OpenStudy (anonymous):

Directions: Type the correct answer in the box. Use numerals instead of words. Neil took his family of five out to dinner at HogTastic Pit BBQ. They all ordered the buffet and drink at $12.80 each. Neil left a 15% gratuity for the waitress on the after-tax bill. Based on a tax rate of 7.5%, Neil ended up paying a total of ___$ for dinner.

OpenStudy (anonymous):

@kohai

OpenStudy (kohai):

... I'm awful at taxes ;-; I have to go to bed, but there should be some people milling around to help you with this

OpenStudy (anonymous):

Okay, thanks anyways.. @kohai

OpenStudy (anonymous):

I'll help you if i can, give me a mo?

OpenStudy (anonymous):

Okay awesome! ^^ a medal? not until you help :\ @ardentpikachu

OpenStudy (anonymous):

I'm not doing it for the medal. :)

OpenStudy (anonymous):

o? :B @ardentpikachu

OpenStudy (anonymous):

I think it goes something like this. Feel free to correct me:

OpenStudy (snowsurf):

The 15% and 7.5% tax rate will be part of the buffet and drink. Total = 12.80 + (12.80)(0.15) + (12.80)(0.075)

OpenStudy (anonymous):

Wait what? so he payed how much? 12.80?? @snowsurf

OpenStudy (snowsurf):

Total = 12.80 + (12.80)(0.15) + (12.80)(0.075) Total = 15.68

OpenStudy (anonymous):

Wrong.. DX @snowsurf

OpenStudy (anonymous):

I think he paid $78.40 altogether. His family of five each ate dinner at $12.80 each 12.80*5=64 He left a 15% gratuity which would be $9.60 He also had a tax rate of 7.5% which would make up $68.80 68.80+9.60= 78.40

OpenStudy (anonymous):

$79.12

OpenStudy (anonymous):

7.5 percent of 64 is 4.8 15 percent of 64 is 9.6 9.6+4.8+64= 78.4

OpenStudy (anonymous):

@gardenGnostic

OpenStudy (anonymous):

Nevermind. ahahaha I forgot that he left the tip AFTER the tax was added onto the bill